At what distance do two beads tighten with a force of 9.8 ∙ 10–17 N if the mass of each bead is 4.5 g?
1
Expert's answer
2021-04-23T11:07:39-0400
"F=G\frac{m^2}{r^2}\\(9.8\cdot10^{-17})=(6.67\cdot10^{-11})\frac{(0.0045)^2}{r^2}\\r=3.7\ m"